lktts:

In the rhombus m able 1 =160. What are m angle 2 and m angle 3?

Vocaloid:

 in a rhombus, which is a type of parallelogram, opposite angles are equal and adjacent angles are supplementary see what you can derive from this

jhonyy9:



jhonyy9:

questions ?

Narad:

Let a and b be the lengths of the diagonals, then \[a^{2}+b^{2}=4*160^{2}\] and the angles will depend on a and b \[\alpha=2*(\cos^{-1} (b/(2+160)))\] and\[\beta=2*(\cos^{-1} (a/(2*160)))\]